OpenStudy (anonymous):

@Mathematics- solve the system by substitution. 3x+7y=36 5x-7=-108 What is the solution of the system?

OpenStudy (turingtest):

are we missing a "y" in the second equation?

OpenStudy (anonymous):

ya. sorry

OpenStudy (saifoo.khan):

x = -12 y = 48/7

OpenStudy (turingtest):

so is it 3x+7y=36 5x-7y=-108 ??????

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you guys give the work

OpenStudy (turingtest):

3x+7y=36 (1) 5x-7y=-108 (2) First solve one equation for y or x (it doesn't matter which) Equation (1) gives 3x=36-7y x=(1/3)(36-7y)=12-7y/3 substitute this expression for x into the other equation, equation (2): 5(12-7y/3)-7y=-108 60-35y/3-21y/3=-108 -56y/3=-168 y=9 substitute this value of y into our expression above for x: x=(1/3)[36-7(9)]=12-21=-9 so we have x=-9, y=9 check by plugging these values into the original equations.

OpenStudy (anonymous):

oh no! you guys all have different answers!

OpenStudy (turingtest):

There was just two of us and saifoo did not see the equation as it was supposed to be (with the missing y) like I said check by plugging these values into the original equation: 3(-9)+7(9)=??? 5(-9)-7(9)=??? use your calcluator (or brain) and you will see that you get the original values: 3(-9)+7(9)=36 5(-9)-7(9)=-108

OpenStudy (turingtest):

and why do you say that?

OpenStudy (turingtest):

@lilg132 Dude read the posts before doing a whole problem. There was a missing "y" in the second equation. You should have solved 3x+7y=36 5x-7y=-108 Please read others work before posting

OpenStudy (lilg132):

kyanmommy can you rewrite the correct question please?

OpenStudy (anonymous):

@Mathematics- solve the system by substitution. 3x+7y=36 5x-7y=-108 What is the solution of the system?

OpenStudy (anonymous):

sorry guys!

OpenStudy (turingtest):

yes, that is mine x=-9 y=9 like I said, check by plugging in

OpenStudy (turingtest):

if you plug in lilg132's answers you get 5(-20.2)-7(13.8)=197.6 so you can see it is wrong.

OpenStudy (turingtest):

sorry 5(-20.2)-7(13.8)=-197.6 is what the second equation becomes with those answers. You should have -108 as you do originally.

OpenStudy (anonymous):

okay im confused. what is the answer

OpenStudy (turingtest):

x=-9 y=9

OpenStudy (anonymous):

okay thanks. Math is not my subject and i am failing thanks for helping

OpenStudy (turingtest):

it's ok, I just wish lilg would concede that he made a mistake and not confuse you.

OpenStudy (anonymous):

lol